# ClearDepth: Enhanced Stereo Perception of Transparent Objects for Robotic Manipulation

### v1

@@ Abstract
Transparent object depth perception poses a challenge in everyday life and logistics, primarily due to the inability of standard 3D sensors to accurately capture depth on transparent or reflective surfaces. This limitation significantly affects depth map and point cloud-reliant applications, especially in robotic manipulation. We developed a vision transformer-based algorithm for stereo depth recovery of transparent objects. This approach is complemented by an innovative feature post-fusion module, which enhances the accuracy of depth recovery by structural features in images. To address the high costs associated with dataset collection for stereo camera-based perception of transparent objects, our method incorporates a parameter-aligned, domain-adaptive, and physically realistic Sim2Real simulation for efficient data generation, accelerated by AI algorithm. Our experimental results demonstrate the model’s exceptional Sim2Real generalizability in real-world scenarios, enabling precise depth mapping of transparent objects to assist in robotic manipulation. Project details are available at https://sites.google.com/view/cleardepth/.
@@ I INTRODUCTION
Transparent objects, such as glass bottles or cups, are commonly found in domestic service robotics and logistics goods sorting scenarios. Their transparent nature, particularly the complexity of refraction and reflection, poses challenges for imaging and recognition [1]. Accurate handling of these objects is crucial for robots that rely on precise three-dimensional information. Whether it’s accurately grasping glass items in home services or efficiently sorting transparent goods at logistics centers, advanced algorithms and technologies are required for machine vision systems. This includes the development of algorithms capable of handling materials with high reflectivity and refraction, or new sensor technologies to better capture the characteristics of transparent objects. These advancements will enhance the ability of robots to handle transparent objects in various environments, push the application of robotic vision technology into broader fields, and enable robots to perform complex tasks more autonomously.
Figure 1: In this work, we explore how to reconstruct depth maps of transparent objects using Sim2Real technology combined with stereo vision, thereby enabling precise manipulation of transparent objects by robots. In the Section III, we focus on introducing a simulation data generation method enhanced by AI, aimed at achieving an optimal balance between the effectiveness and speed of data generation. The Section IV conducts both qualitative and quantitative analyses comparing the performance of our method with the baseline, and demonstrates the efficacy of our approach through robot grasping experiments.
@@
In addressing the depth perception challenges posed by transparent objects, researchers have developed innovative approaches. Initial efforts focused on using deep learning techniques to reconstruct depth information from incomplete depth images captured by RealSense cameras, optimizing for greater accuracy [2, 3]. Subsequent studies explored stereo vision systems [4] and multi-view approaches [5], leveraging neural networks to enhance perception capabilities. Despite these advancements, real-world applications still face challenges, such as inconsistent depth inputs and the complexities of multi-view imaging. A major obstacle in 3D reconstruction of transparent objects is the scarcity of reliable feature points. Innovations addressing this include extracting structural details, such as boundary identification [6], and improving hardware capabilities for more accurate depth imagery [7]. Deep learning has played a pivotal role in learning and interpreting the complex geometric features of these objects, thereby enhancing the precision and reliability of 3D models. For deep learning tasks involving transparent objects, extensive and accurate datasets are crucial. Various techniques have been used for dataset collection, including using markers for object poses [8, 9], substituting transparent objects with opaque ones [10], and manual manipulation in 3D software [11]. However, these methods are labor-intensive and often result in noisy depth maps. To overcome these challenges, simulation engines are increasingly used to generate synthetic data [11, 5], though balancing the realism of ray-tracing with the efficiency of rasterization remains a significant challenge. Our approach includes developing an AI-accelerated tool for generating realistic simulation datasets for transparent object depth recovery. This tool enables direct application of models on actual sensors without post-training, producing instance segmentation, poses, and depth maps. We also introduce a stereo depth recovery network specifically designed for transparent objects. Utilizing a cascaded vision transformer (ViT) backbone, it efficiently extracts contextual structural information. This network outperforms previous CNN and ViT-based models, particularly for high-resolution images and robotic manipulation. Our contributions address the limitations of existing depth estimation methods and datasets for transparent objects. We present a novel AI-enhanced stereo-based simulation dataset generation method, creating a detailed, noise-free dataset. Additionally, our end-to-end stereo depth recovery network operates independently of mask priors and includes a post-processing structure that enhances structural details. This stereo imaging algorithm can be directly integrated into existing robotic grasping pipelines. The whole pipeline is as in Fig. 1.
@@ II-A Perception of Transparent Objects
Robotics faces significant challenges in perceiving transparent objects due to their low contrast and complex light behavior, which affects sensor accuracy in determining their position and shape. Traditional sensors like RGB and RGB-D cameras often struggle with transparent objects because they focus on intensity data, missing the nuances of optical properties. To address this, research has increasingly turned to polarized cameras, which reduce reflections, increase contrast, and provide richer light information, enhancing the detection of transparent objects’ shape and position [7, 12]. However, their high cost limits widespread use in robotics. Some studies, like [13], enhance transparent object tracking by combining CNN and transformer-based encoders, while [14] employs Sim2Real strategies with simulated datasets. Research such as [15] addresses transparent object matting by predicting alpha channel values, aiding image editing. For robotic manipulation and pose estimation, recent works have expanded perception models to include multi-task extensions [8, 11, 16]. Depth recovery and reconstruction remain particularly challenging due to light interactions, with techniques like Nerf and volumetric rendering used for surface reconstruction [17, 7, 18, 19], and stereo and multiview approaches for depth and geometry regression [4, 20, 9]. These efforts utilize various sensors, including RGB-D, stereo-vision, and multi-view systems, to improve transparent object perception [3, 2, 8, 11, 16, 9, 4, 5, 18, 7]. The field, while complex, is advancing with deep learning and sensor technology, aiming to enhance accuracy and reliability.
@@ II-B Deep Learning-based Stereo Depth Recovery
Deep learning-based stereo matching methods have recently outperformed traditional approaches, with 2D convolutional models [21, 22] offering simplicity and efficiency. These models achieve high accuracy even on limited computational resources, making them suitable for engineering applications, though they still require improvements in accuracy and robustness due to 3D cost space constraints. 3D convolutional networks [23, 24] provide better interpretability and higher disparity map accuracy but require optimization due to their computational demands. RAFT-Stereo [25], an extension of RAFT [26], applies optical flow techniques to stereo matching, improving generalization and robustness with its lightweight GRU module, but struggles with global context extraction due to its CNN architecture. STTR [27], inspired by SuperGlue [28], uses transformers with positional embedding and attention mechanisms for binocular dense matching, producing disparity and depth maps. However, these methods are computationally intensive and slow in inference, limiting their suitability for high-resolution images and downstream robotic tasks.
@@ II-C Transparent Object Datasets
In the open-source community, various datasets for transparent objects are available, such as real-world datasets [8, 29, 30]. In Depth Anything v1, the authors used both open-source real and synthetic datasets for training, but in v2, they switched entirely to synthetic data. Introducing even 10% real data can significantly degrade performance due to sensor noise in the labels, which is particularly detrimental in depth recovery tasks. The challenge of accurately annotating complex, occluded real-world scenes led us to prioritize synthetic data. To mitigate potential sim2real gaps, we used ray-tracing renderers to simulate realistic light refraction. Existing synthetic datasets, such as  [10, 31, 9, 32, 33, 16], typically feature transparent objects on desktops. However, these datasets lack the complexity necessary for generalization to real-world scenarios like kitchens, bedrooms, and offices, where service and humanoid robots operate. Moreover, these datasets often require extensive pre- or post-processing, such as segmentation or background reconstruction, which is impractical for end-to-end algorithms crucial to embodied intelligence applications. Synthetic datasets with HDRI backgrounds, such as [5], also face challenges. HDRI backgrounds lack depth value labels, leading to confusion and poor generalization, especially in zero-shot scenarios for service robots. Simulation data for Realsense cameras, like  [16, 33], involve complex rendering processes that reduce efficiency. The process includes rendering the RGB image, simulating IR projection with spotlights, and generating noisy depth maps, which is both complex and time-consuming. In designing our dataset and generation process, we address the limitations of existing datasets while anticipating the needs of future embodied intelligence algorithms. Our dataset includes indoor scenes typical for service robots, enriched with background depth values for transparent objects. It features both container-type and cosmetic transparent objects, carefully designed to overcome current deficiencies and ensure future scalability.
@@ III-A Network Overview
Figure 2: Our stereo depth recovery network for transparent objects. Feature extraction is performed on left and right images, with additional processing by a shared weighted context encoder for the left image. A correlation pyramid, created by merging the feature maps, is refined through a GRU loop, enhancing structural information for depth recovery. The output is a refined disparity map.
Transparent objects refract background textures, making texture features less relevant and structural features more critical for imaging and perception tasks. In [34], it was found that CNNs excel at recognizing textures, while ViTs are better at recognizing shapes. Traditional ViTs downsample inputs and outputs to a smaller size, using learnable upsampling. While effective in many vision tasks, ViTs are challenging to train, computationally intensive, and less effective at fine-grained feature extraction. To address this, models like SegFormer [35] and DinoV2 [36] use cascaded structures to enhance ViTs for fine-grained tasks like depth estimation and semantic segmentation, providing multi-scale features and easing training.
For transparent objects, structural features are crucial. Considering the fine-grained needs of stereo tasks and the demand for lightweight, fast models in robotics, we utilize MixVisionTransformer B5 for feature extraction. Traditional stereo deep learning networks rely on feature dot products for similarity, which is inadequate for transparent objects due to background texture refraction. Incorporating structural feature priors in the GRU loop is essential. While cross-attention or spatial-attention structures could be used, they are parameter-heavy and slow in inference. Therefore, we designed a lightweight structural feature post-fusion module to introduce these features into the GRU loop, enhancing performance by leveraging the low-level structural feature. Our network structure is as in Fig. 2.
@@
Our backbone architecture, conforming to the B5 settings of SegFormer, initiates with overlap patch embedding for initial image tokenization, crucial for preserving local features. This process sequentially passes the tokens through four transformer blocks ($N=4$), effectively generating feature maps at scaled dimensions of $\frac{1}{4}$, $\frac{1}{8}$, $\frac{1}{16}$, and $\frac{1}{32}$. To optimize computational efficiency, the model incorporates efficient self-attention, which significantly reduces the computational burden from $O(N^{2})$ to $O(\frac{N^{2}}{R})$ by implementing a reduction ratio $R$. This reduction is achieved by first reshaping the input sequence from $N\cdot C$ to $\frac{N}{R}\times(C\cdot R)$ by 2d convolutional layer with the stride $8,4,2,1$ for different ViT blocks, as detailed in equation 1, and then adjusting the sequence dimensions back to $C$ through linear layers, as described in equation 2.
@@
Figure 3: We utilize hardware-accelerated ray tracing, OptiX AI denoiser, and AI-driven super-resolution. This process accelerates the dataset generation process.
@@ III-A2 Structural Feature Post-Fusion
Here, $x_{k}$ is a concatenation of several feature maps, including the correlation $\mathbf{C}_{k}$, the current disparity $\mathbf{d}_{k}$, and structural context features $\mathbf{c}_{k}$, $\mathbf{c}_{r}$, and $\mathbf{c}_{h}$. Specifically, $\mathbf{c}_{k}$, $\mathbf{c}_{r}$, and $\mathbf{c}_{h}$ represent structural features derived from the left image. These features are incorporated as residuals into the GRU loop, allowing for enhanced participation of structural information during the disparity map refinement process.
@@
In summary, to address the challenges of transparent objects, we selected an appropriate image feature extractor. Additionally, considering the unique difficulties of transparent objects and the need for lightweight models in robotics, we designed a structural feature post-fusion architecture. Every detail of our network structure is tailored to the characteristics of transparent object scenarios.
In the comparative experiments section, the visual results demonstrate that our model significantly enhances the stereo imaging of transparent objects.
@@ III-B Synthetic Dataset Generation
Figure 5: Sample images from our SynClearDepth dataset are presented, showcasing randomly placed transparent objects in various indoor scenes (bathroom, dining room, kitchen, living room) under different lighting conditions. The objects include common transparent items like cosmetic packaging, glass containers, etc.
To enhance dataset generation efficiency, we utilized OptiX’s AI denoiser and deep learning super-resolution techniques [37], notably reducing the time from an average of $12.77$ seconds to $4.40$ seconds per set (including stereo RGB, depth, masks, and object-camera poses), significantly speeding up the process by cutting nearly two-thirds data generation time while maintaining high-quality outputs. Our dataset generation process is illustrated in Fig. 3. Our proposed dataset SynClearDepth comprises 16 selected objects, including 10 common transparent containers in daily life and 6 products with glass material, as shown in Fig. 4. To ensure that the background of the generated dataset also contains depth value and avoid ambiguities in depth recovery training due to lack of dataset labeling of background part, we collected indoor scenes in combination with the object models, encompassing 6 bathrooms, 3 dining rooms, 5 kitchens, and 6 living rooms. Through this combination, we generated 14,091 image sets, each containing left and right RGB images with size 1280*720, ground truth depth, instance segmentation maps, as well as object and camera poses, as illustrated in Fig. 5. In the dataset generation process, we applied domain randomization to object types, quantities, positions and poses, lighting conditions, and camera shooting angles.
III-C metrics
@@
IV Experiments
IV-A Technical Specifications
IV-B Qualitative and Quantitative Studies for Stereo Depth Estimation
IV-B1 Quantitative Analysis on Middlebury Dataset
@@
TABLE I: Quantitative results on Middleburry Stereo Evaluation Benchmark [41]. All metrics have been calculated using undisclosed weighting factors. The outcomes unequivocally demonstrate that our technique significantly outperforms the baseline method.
IV-B2 Quantitative Analysis on KITTI Dataset
((a)) Left image
((b)) RAFT-Stereo[25]
((c)) StereoBase[43]
((d)) Ours
IV-B3 Evaluation on Our Transparent Object Dataset
@@
clearDepth 2.138 8.7282 24.7329 16.3178 9.8459 5.76
TABLE II: Quantitative results on transparent object dataset compared with stereo SOTA methods.

### v2

# ClearDepth: Efficient Stereo Perception of Transparent Objects for Robotic Manipulation

@@ Abstract
Transparent object depth perception remains a major challenge in robotics and logistics due to the limitations of standard 3D sensors in capturing accurate depth on transparent and reflective surfaces. This affects applications relying on depth maps and point clouds, particularly in robotic manipulation. To address this, we propose a vision transformer-based algorithm for stereo depth recovery of transparent objects, enhanced by a novel feature post-fusion module that refines depth estimation using structural image features. To mitigate the high costs of stereo dataset collection, we introduce a physically realistic, domain-adaptive Sim2Real framework for efficient data generation. Our method, accelerated by AI, demonstrates strong Sim2Real generalization, enabling precise depth mapping of transparent objects for robotic applications. Dataset and project details are available at https://sites.google.com/view/cleardepth/.
@@ I INTRODUCTION
Transparent objects, such as glass bottles or cups, are commonly found in domestic service robotics and logistics goods sorting scenarios. Their transparent nature, particularly the complexity of refraction and reflection, poses challenges for imaging and recognition [1]. Accurate handling of these objects is crucial for robots that rely on precise three-dimensional information. Whether it’s accurately grasping glass items in home services or efficiently sorting transparent goods at logistics centers, advanced algorithms and technologies are required for machine vision systems. This includes the development of algorithms capable of handling materials with high reflectivity and refraction, or new sensor technologies to better capture the characteristics of transparent objects. These advancements will enhance the ability of robots to handle transparent objects in various environments, push the application of robotic vision technology into broader fields, and enable robots to perform complex tasks more autonomously. In addressing the depth perception challenges posed by transparent objects, researchers have developed innovative approaches. Initial efforts focused on using deep learning techniques to reconstruct depth information from incomplete depth images captured by RealSense cameras, optimizing for greater accuracy [2, 3]. Subsequent studies explored stereo vision systems [4] and multi-view approaches [5], leveraging neural networks to enhance perception capabilities. Despite these advancements, real-world applications still face challenges, such as inconsistent depth inputs and the complexities of multi-view imaging. A major obstacle in 3D reconstruction of transparent objects is the scarcity of reliable feature points. Innovations addressing this include extracting structural details, such as boundary identification [6], and improving hardware capabilities for more accurate depth imagery [7]. Deep learning has played a pivotal role in learning and interpreting the complex geometric features of these objects, thereby enhancing the precision and reliability of 3D models. For deep learning tasks involving transparent objects, extensive and accurate datasets are crucial. Various techniques have been used for dataset collection, including using markers for object poses [8, 9], substituting transparent objects with opaque ones [10], and manual manipulation in 3D software [11]. However, these methods are labor-intensive and often result in noisy depth maps. To overcome these challenges, simulation engines are increasingly used to generate synthetic data [11, 5], though balancing the realism of ray-tracing with the efficiency of rasterization remains a significant challenge. Our approach includes developing an AI-accelerated tool for generating realistic simulation datasets for transparent object depth recovery. This tool enables direct application of models on actual sensors without post-training, producing instance segmentation, poses, and depth maps. We also introduce a stereo depth recovery network specifically designed for transparent objects. Utilizing a cascaded vision transformer (ViT) backbone, it efficiently extracts contextual structural information. This network outperforms previous CNN and ViT-based models, particularly for high-resolution images and robotic manipulation. Our contributions address the limitations of existing depth estimation methods and datasets for transparent objects. We present a novel AI-enhanced stereo-based simulation dataset generation method, creating a detailed, noise-free dataset. Additionally, our end-to-end stereo depth recovery network operates independently of mask priors and includes a post-processing structure that enhances structural details. This stereo imaging algorithm can be directly integrated into existing robotic grasping pipelines. The whole pipeline is shown in Fig. LABEL:fig.cleardepth_overview.
@@ II-A Perception of Transparent Objects
Robotic perception of transparent objects remains challenging due to their low contrast and complex light interactions, which affect sensor accuracy in determining position and shape. Traditional RGB and RGB-D cameras struggle with these objects, as they rely on intensity data and overlook optical properties. To address this, research has explored polarized cameras, which reduce reflections and enhance contrast [7, 12]. However, their high cost limits widespread adoption. Alternative approaches include CNN-transformer-based models for tracking [13], Sim2Real techniques leveraging synthetic datasets [14], and alpha-matting methods for transparent object segmentation [15]. For robotic manipulation and pose estimation, multi-task perception models have been introduced [8, 11, 16]. Depth recovery remains particularly challenging due to light refraction and reflection. Methods such as NeRF and volumetric rendering aid surface reconstruction [17, 18, 19], while stereo and multi-view techniques improve depth estimation [4, 20, 9]. These approaches leverage various sensors, including RGB-D, stereo vision, and multi-view systems, to enhance transparent object perception [3, 2, 8, 11, 16, 9, 4, 5, 18, 7]. Advances in deep learning and sensor technologies continue to drive improvements in accuracy and reliability.
@@ II-B Deep Learning-based Stereo Depth Recovery
Deep learning-based stereo matching methods have recently outperformed traditional approaches, with 2D convolutional models [21, 22] offering simplicity and efficiency. These models achieve high accuracy even on limited computational resources, making them suitable for engineering applications, though they still require improvements in accuracy and robustness due to 3D cost space constraints. 3D convolutional networks [23, 24] provide better interpretability and higher disparity map accuracy but require optimization due to their computational demands. RAFT-Stereo [25], an extension of RAFT [26], applies optical flow techniques to stereo matching, improving generalization and robustness with its lightweight Gated Recurrent Unit (GRU) Network module, but struggles with global context extraction due to its CNN architecture. STTR [27], inspired by SuperGlue [28], uses transformers with positional embedding and attention mechanisms for binocular dense matching, producing disparity and depth maps. However, these methods are computationally intensive and slow in inference, limiting their suitability for high-resolution images and downstream robotic tasks.
@@ II-C Transparent Object Datasets
Recent works [29, 30] show that real-world datasets degrade model performance due to label noise, while synthetic data with precise labels, enhance model performance. However, synthetic datasets across different data domains remain scarce in the open-source community. Most recent vision tasks leverage ray-tracing renderers, minimizing the sim2real gap, making data domain gaps the primary factor affecting model performance. Therefore, continuously expanding synthetic datasets in the open-source community is crucial. Existing synthetic datasets, such as  [10, 31, 9, 32, 33, 16], typically feature transparent objects on desktops. However, these datasets lack the complexity necessary for generalization to real-world scenarios like kitchens, bedrooms, and offices, where service and humanoid robots operate. Moreover, these datasets often require extensive pre- or post-processing, such as segmentation or background reconstruction, which is impractical for end-to-end algorithms crucial to embodied intelligence applications. Synthetic datasets with HDRI backgrounds, such as [5], also face challenges. HDRI backgrounds lack depth value labels, leading to confusion and poor generalization, especially in zero-shot scenarios for service robots. Simulation data for Realsense cameras, like  [16, 33], involve complex rendering processes that reduce efficiency. The process includes rendering the RGB image, simulating IR projection with spotlights, and generating noisy depth maps, which is both complex and time-consuming. In designing our dataset and generation process, we address the limitations of existing datasets while anticipating the needs of future embodied intelligence algorithms. Our dataset includes indoor scenes typical for service robots, enriched with background depth values for transparent objects. It features both container-type and cosmetic transparent objects, carefully designed to overcome current deficiencies and ensure future scalability.
@@ III-A Network Overview
Figure 2: Our stereo depth recovery network for transparent objects. Feature extraction is performed on left and right images, with additional processing by a shared weighted context encoder for the left image. A correlation pyramid, created by merging the feature maps, is refined through a GRU loop, enhancing structural information for depth recovery. The output is a refined disparity map. The feature encoder is used for computing visual similarity for left and right image, the context encoder extracts features for disparity refinement in the GRU.
Transparent objects refract background textures, making structural features more critical than texture features for imaging and perception. [34] shows CNNs excel at texture recognition, while ViTs are better at shape recognition. Traditional ViTs downsample inputs and use learnable upsampling, which, while effective, is computationally costly and weak in fine-grained feature extraction. Models like SegFormer [35] and DinoV2 [36] enhance ViTs with cascaded structures, improving depth estimation and segmentation through multi-scale features. To better capture structural features for transparent objects, we adopt an efficient cascaded ViT backbone. Traditional stereo networks rely on dot product similarity, which fails for transparent objects due to background refraction. Instead, we introduce structural feature priors into the GRU loop via a lightweight post-fusion module, enhancing performance without the overhead of cross-attention. To ensure adaptability to robotic tasks, we keep the model architecture lightweight to avoid potential engineering challenges. Our network is illustrated in Fig. 2.
@@
Our backbone begins with overlap patch embedding for initial tokenization, preserving local features. Tokens pass through four transformer blocks, generating feature maps at $\frac{1}{4}$, $\frac{1}{8}$, $\frac{1}{16}$, and $\frac{1}{32}$ scales. To optimize computational efficiency, the model incorporates efficient self-attention, which significantly reduces the computational burden from $O(N^{2})$ to $O(\frac{N^{2}}{R})$. This reduction is achieved by first reshaping the input sequence from $N\cdot C$ to $\frac{N}{R}\times(C\cdot R)$ by 2d convolutional layer with the stride $8,4,2,1$ for different ViT blocks, as detailed in Equ. 1, and then adjusting the sequence dimensions back to $C$ channel through linear layers, as described in Equ. 2. K denotes the sequence in the ViT block optimized for lower computational complexity.
@@ III-A2 Structural Feature Post-Fusion
Here, $x_{k}$ is a concatenation of several feature maps, including the correlation $\mathbf{C}_{k}$, the current disparity $\mathbf{d}_{k}$, and structural context feature maps $\mathbf{c}_{k}$, $\mathbf{c}_{r}$, and $\mathbf{c}_{h}$. Specifically, $\mathbf{c}_{k}$, $\mathbf{c}_{r}$, and $\mathbf{c}_{h}$ represent structural features derived from the left image. These features are incorporated as residuals into the GRU loop, allowing for enhanced participation of structural information during the disparity map refinement process. $z,r,h$ represent the state information of the update gate, reset gate, and hidden gate in a GRU.
Figure 3: We utilize hardware-accelerated ray tracing, pretrained OptiX AI denoiser, and pretrained AI-driven super-resolution. This process accelerates the dataset generation process.
@@
In summary, to address the challenges of transparent objects, we selected an appropriate image feature extractor. Additionally, considering the unique difficulties of transparent objects and the need for efficient models in robotics, we designed a structural feature post-fusion architecture. Every detail of our network structure is tailored to the characteristics of transparent object scenarios.
In the comparative experiments section, the visual results demonstrate that our model substantially enhances the stereo imaging of transparent objects.
@@ III-B Synthetic Dataset Generation
Figure 5: Sample images from SynClearDepth dataset, depicting transparent objects randomly placed in indoor scenes (bathroom, dining room, kitchen, living room) under various lighting conditions. The objects, including cosmetic packaging and glass containers, are randomly dropped onto tables using 3D bounding boxes as collision bodies.
To enhance the efficiency of synthetic dataset generation, we utilized the AI denoiser provided by OptiX [37] during rendering and adopted open-source pretrained deep learning super-resolution [38] as rendering output optimization strategies, reducing the average generation time per set (stereo RGB, depth, masks, and object-camera poses) from $12.77$ to $4.40$ seconds. Since these techniques are widely used in computer graphics and do not alter the core data distribution, we omit further analysis. The dataset generation process is illustrated in Fig. 3. Our SynClearDepth dataset includes 16 selected objects: 10 common transparent containers and 6 glass-material products (Fig. 4). To ensure depth labels for both objects and backgrounds, we combined object models with indoor scenes, including 6 bathrooms, 3 dining rooms, 5 kitchens, and 6 living rooms. This resulted in 14,091 image sets, each containing left and right RGB images, ground truth depth, instance segmentation, and object/camera poses (Fig. 5). We applied domain randomization to object types, quantities, poses, lighting, and camera angles. This dataset is designed to support robotic perception and manipulation in service robot applications, particularly for handling transparent objects in household environments.
IV Experiments
IV-A Technical Specifications
IV-B Evaluation Metrics
@@
TABLE I: Quantitative results on Middleburry Stereo Evaluation Benchmark [42]. All metrics have been calculated using undisclosed weighting factors. The outcomes unequivocally demonstrate that our technique substantially outperforms the baseline method.
@@
ClearDepth 2.138 8.7282 24.7329 16.3178 9.8459 5.7600
TABLE II: Quantitative results on transparent object dataset compared with stereo SOTA methods. Visualization results shown in Fig. 7. Our efficient stereo depth estimation model outperforms other SOTA models in transparent object dataset whtile maintaining lower computational cost. Results show that our model achieves strong performance in average numerical error and outperforms all others in the proportion of error pixels.
IV-C Qualitative and Quantitative Studies for Stereo Depth Estimation
IV-C1 Quantitative Analysis on Middlebury Dataset
IV-C2 Quantitative Analysis on KITTI Dataset
(a) Left image
(b) RAFT-Stereo[25]
(c) StereoBase[46]
(d) Ours
IV-C3 Evaluation on Transparent Object Dataset